If m is divisible by 5, what is the largest number that must be evenly divisible into 5m+25?
Answer by stanbon(75887)

If m is divisible by 5, what is the largest number that must be evenly divisible into 5m+25?
----
If m is divisible by 5, m = 5k where k is an integer NOT divisible
by 5.
---
So 5m+25 = 5(5k)+25 = 25(k+1)
Therefore 5m+25 is divisible by 25
=============
But that is not the end of the story because
k+1 may also be divisble by 5.
================
Cheers,
Stan H.